Tom is 3 years older than Esmeralda. In 6 years the sum of their ages will be 65. How old is Tom now?
Answer by ankor@dixie-net.com(22740)

let t = Tom's age now
let e = Esmeralda's age now
:
Write an equation for each statement
:
Tom is 3 years older than Esmeralda.
t = e + 3
or
e = t - 3
:
In 6 years the sum of their ages will be 65.
(t+6) + (e+6) = 65
t + e + 12 = 65
t + e = 65 - 12
t + e = 53
How old is Tom now?
Replace e with (t-3) in the above equation
t + (t-3) = 53
2t = 53 + 3
2t = 56
t = 56/2
t = 28 yrs is Tom's present age
:
:
See if that checks out, e=25
28 + 6 + 25 + 6 = 65